OVHCloud Bare Metal Cloud Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#3944 Switch replacements in rooms 1, 2 and 3
Scheduled Maintenance Report for Bare Metal Cloud
Completed
We start a series of switches replacement in our data center in Paris 19.
The entire park will be replaced by newer models with new features.

The replacement will be performed at night, causing a break of 2-5 minutes on each server.




Update(s):

Date: 2010-05-11 14:03:48 UTC
This morning, we are carrying out the replacement of the switch of the A01 bay.
The next ones will be done at night.
Posted Apr 27, 2010 - 14:32 UTC